# On the Classification of Two-Dimensional Algebras

### 16 Mar 2026 v1

@@ 1. Introduction
Our original aim was to derive this number using the classification result [3]. However, the difference between the above formula and the number obtained from the classification result forced us to recheck the classification one more time. It turned out that although the listed canonical classes of algebras are correct, the description of the isomorphisms between elements within two of these classes contains certain inaccuracies.
In this paper we first correct these inaccuracies and then use the corrected classification to determine the number of non-isomorphic two-dimensional algebras over a finite field.
@@ 2. On the Classification of Two-Dimensional Algebras over an Arbitrary Base Field
If $\eta_{1}=0$, then $\alpha^{\prime}_{2}=\alpha^{\prime}_{4}=0$ and $\alpha^{\prime}_{1}=\xi_{1}\alpha_{1},\beta^{\prime}_{1}=\frac{\xi^{2}_{1}}{\eta_{2}}\left(\beta_{1}-3\alpha_{1}\frac{\xi_{2}}{\xi_{1}}\right).$ Therefore, if $\alpha_{1}\neq 0$, we may assume $\alpha^{\prime}_{1}=1,\beta^{\prime}_{1}=0$ to get $A^{\prime}=\left(\begin{array}[]{cccc}1&0&0&0\\ 0&-1&-1&0\end{array}\right)$, which is isomorphic to $A_{12}(0),$ if $\alpha_{1}=0$, then $\alpha^{\prime}_{1}=0$ and we may assume $\beta^{\prime}_{1}=1$ to get $A_{13}=\left(\begin{array}[]{cccc}0&0&0&0\\ 1&0&0&0\end{array}\right).$
@@
Case 2-4: $\alpha_{1}=0,\alpha_{2}=0$. In this case one has $A_{12,2}=\left(\begin{array}[]{cccc}0&0&0&0\\ 1&0&0&0\end{array}\right).$
@@
If $\eta_{1}=0$, then $\alpha^{\prime}_{2}=\alpha^{\prime}_{4}=0$ and $\alpha^{\prime}_{1}=\xi_{1}\alpha_{1},\beta^{\prime}_{1}=\frac{\xi^{2}_{1}}{\eta_{2}}\beta_{1}.$ Therefore, one can have $A^{\prime}=\left(\begin{array}[]{cccc}1&0&0&0\\ 0&-1&-1&0\end{array}\right)$ or $A_{12,3}=\left(\begin{array}[]{cccc}1&0&0&0\\ 1&-1&-1&0\end{array}\right)$ or $A_{13,3}=\left(\begin{array}[]{cccc}0&0&0&0\\ 1&0&0&0\end{array}\right)$. In the first case $A^{\prime}$ is isomorphic to $A_{12,3}(0).$
Conclusion. In the Theorem 1.1 the items $A_{10},A_{11}$, $A_{8,2},A_{9,2}$, $A_{10,3},A_{11,3}$ should be replaced by the following items, respectively:
@@ 3. On the Number of Non-Isomorphic Two-Dimensional Algebras over a Finite Field
  • $|A_{1}|=q^{4},|A_{2}|=q^{2}(q-1),|A_{3}|=3q^{2},|A_{4}|=q^{2},|A_{5}|=q,|A_{6}|=q(q-1),|A_{7}|=3q,|A_{8}|=q,|A_{9}|=1,|A_{11}|=\left\{\begin{array}[]{ll}2&,\,\,if\,\,3\,\,divides\,\,q-2,\\ 4&,\,\,otherwise.\end{array}\right.,|A_{12}|=3,|A_{13}|=1$ and the total number of them, including the trivial algebra, is
    $\left\{\begin{array}[]{ll}q^{4}+q^{3}+4q^{2}+4q+8&,\,\,if\,\,3\,\,divides\,\,q-2,\\ q^{4}+q^{3}+4q^{2}+4q+10&,\,\,otherwise.\end{array}\right$.
  • $|A_{1,2}|=q^{4},|A_{2,2}|=q^{2}(q-1)+q,|A_{3,2}|=2q^{2},|A_{4,2}|=2q^{2}-q,|A_{5,2}|=q(q-1)+1,|A_{6,2}|=2q,|A_{7,2}|=2q-1,|A_{9,2}|=\left\{\begin{array}[]{ll}2&,\,if\,q=2^{2k-1},\\ 4&,\,if\,q=2^{2k}\end{array}\right.,|A_{10,2}|=2,|A_{11,2}|=1,\\ |A_{12,2}|=1$ and the total number of them, including the trivial algebra, is
    $\left\{\begin{array}[]{ll}q^{4}+q^{3}+4q^{2}+3q+7&,\,if\,q=2^{2k-1},\\ q^{4}+q^{3}+4q^{2}+3q+8&,\,if\,q=2^{2k}.\end{array}\right$.
  • $|A_{1,3}|=q^{4},|A_{2,3}|=q^{2}(q-1),|A_{3,3}|=3q^{2},|A_{4,3}|=q^{2},|A_{5,3}|=q,|A_{6,3}|=q(q-1),|A_{7,3}|=3q,|A_{8,3}|=q,|A_{10,3}|=2,|A_{11,3}|=3,|A_{12,3}|=1,|A_{13,3}|=1$ and the total number of them, including the trivial algebra, is
    $q^{4}+q^{3}+4q^{2}+4q+8.$
@@
    $\left\{\begin{array}[]{ll}q^{4}+q^{3}+4q^{2}+4q+8+|A_{10}|&,\,\,if\,\,3\,\,divides\,\,q-2,\\ q^{4}+q^{3}+4q^{2}+4q+10+|A_{10}|&,\,\,otherwise.\end{array}\right$.
@@
    $\left\{\begin{array}[]{ll}q^{4}+q^{3}+4q^{2}+3q+7+|A_{8,2}|&,\,if\,q=2^{2k-1},\\ q^{4}+q^{3}+4q^{2}+3q+8+|A_{8,2}|&,\,if\,q=2^{2k}.\end{array}\right$.
@@
Hence these partially defined transformations form a semigroup action on the parameter set and there is a need for the number of orbits of this action.

### 29 Mar 2026 v2

@@ 1. Introduction
Our original aim was to derive this number using the classification result [3]. However, the difference between the above formula and the number obtained from the classification result forced us to recheck the classification one more time. It turns out that some corrections are needed in the classification of the fifth family of two-dimensional algebras in [3].
In this paper, we first provide these corrections and then use the revised classification to determine the number of non-isomorphic two-dimensional algebras over a finite field.
@@ 2. On the Classification of Two-Dimensional Algebras over an Arbitrary Base Field
If $\eta_{1}=0$, then $\alpha^{\prime}_{2}=\alpha^{\prime}_{4}=0$ and $\alpha^{\prime}_{1}=\xi_{1}\alpha_{1},\beta^{\prime}_{1}=\frac{\xi^{2}_{1}}{\eta_{2}}\left(\beta_{1}-3\alpha_{1}\frac{\xi_{2}}{\xi_{1}}\right).$ Therefore, if $\alpha_{1}\neq 0$, we may assume $\alpha^{\prime}_{1}=1,\beta^{\prime}_{1}=0$ to get $A^{\prime}=\left(\begin{array}[]{cccc}1&0&0&0\\ 0&-1&-1&0\end{array}\right)$, which is isomorphic to $A_{12}(0),$ if $\alpha_{1}=0$, then $\alpha^{\prime}_{1}=0$ and we may assume $\beta^{\prime}_{1}=1$ to get $A=\left(\begin{array}[]{cccc}0&0&0&0\\ 1&0&0&0\end{array}\right),$ which is isomorphic to $A_{11}(0).$
@@
Case 2-4: $\alpha_{1}=0,\alpha_{2}=0$. In this case one has $A=\left(\begin{array}[]{cccc}0&0&0&0\\ 1&0&0&0\end{array}\right)\simeq A_{9,2}(0).$
@@
If $\eta_{1}=0$, then $\alpha^{\prime}_{2}=\alpha^{\prime}_{4}=0$ and $\alpha^{\prime}_{1}=\xi_{1}\alpha_{1},\beta^{\prime}_{1}=\frac{\xi^{2}_{1}}{\eta_{2}}\beta_{1}.$ Therefore, one can have $A^{\prime}=\left(\begin{array}[]{cccc}1&0&0&0\\ 0&-1&-1&0\end{array}\right)\simeq A_{11,3}(0)$, or $A_{12,3}=\left(\begin{array}[]{cccc}1&0&0&0\\ 1&-1&-1&0\end{array}\right)\simeq A_{9,3}(0)$ or $A=\left(\begin{array}[]{cccc}0&0&0&0\\ 1&0&0&0\end{array}\right)\simeq A_{10,3}(0)$.
Conclusion. In the Theorem 1.1 the items $A_{13},A_{12,2},A_{12,3},A_{13,3}$ should be omitted and the items $A_{10},A_{11}$, $A_{8,2},A_{9,2}$, $A_{10,3},A_{11,3}$ should be replaced by the following items, respectively:
@@ 3. On the Number of Non-Isomorphic Two-Dimensional Algebras over a Finite Field
  • $|A_{1}|=q^{4},|A_{2}|=q^{2}(q-1),|A_{3}|=3q^{2},|A_{4}|=q^{2},|A_{5}|=q,|A_{6}|=q(q-1),|A_{7}|=3q,|A_{8}|=q,|A_{9}|=1,|A_{11}|=\left\{\begin{array}[]{ll}2&,\,\,if\,\,3\,\,divides\,\,q-2,\\ 4&,\,\,otherwise.\end{array}\right.,|A_{12}|=3$ and the total number of them, including the trivial algebra, is
    $\left\{\begin{array}[]{ll}q^{4}+q^{3}+4q^{2}+4q+7&,\,\,if\,\,3\,\,divides\,\,q-2,\\ q^{4}+q^{3}+4q^{2}+4q+9&,\,\,otherwise.\end{array}\right$.
  • $|A_{1,2}|=q^{4},|A_{2,2}|=q^{2}(q-1)+q,|A_{3,2}|=2q^{2},|A_{4,2}|=2q^{2}-q,|A_{5,2}|=q(q-1)+1,|A_{6,2}|=2q,|A_{7,2}|=2q-1,|A_{9,2}|=\left\{\begin{array}[]{ll}2&,\,if\,q=2^{2k-1},\\ 4&,\,if\,q=2^{2k}\end{array}\right.,|A_{10,2}|=2,|A_{11,2}|=1$ and the total number of them, including the trivial algebra, is
    $\left\{\begin{array}[]{ll}q^{4}+q^{3}+4q^{2}+3q+6&,\,if\,q=2^{2k-1},\\ q^{4}+q^{3}+4q^{2}+3q+7&,\,if\,q=2^{2k}.\end{array}\right$.
  • $|A_{1,3}|=q^{4},|A_{2,3}|=q^{2}(q-1),|A_{3,3}|=3q^{2},|A_{4,3}|=q^{2},|A_{5,3}|=q,|A_{6,3}|=q(q-1),|A_{7,3}|=3q,|A_{8,3}|=q,|A_{10,3}|=2,|A_{11,3}|=3$ and the total number of them, including the trivial algebra, is
    $q^{4}+q^{3}+4q^{2}+4q+6.$
@@
    $\left\{\begin{array}[]{ll}q^{4}+q^{3}+4q^{2}+4q+7+|A_{10}|&,\,\,if\,\,3\,\,divides\,\,q-2,\\ q^{4}+q^{3}+4q^{2}+4q+9+|A_{10}|&,\,\,otherwise.\end{array}\right$.
@@
    $\left\{\begin{array}[]{ll}q^{4}+q^{3}+4q^{2}+3q+6+|A_{8,2}|&,\,if\,q=2^{2k-1},\\ q^{4}+q^{3}+4q^{2}+3q+7+|A_{8,2}|&,\,if\,q=2^{2k}.\end{array}\right$.
@@
Hence there is a need for an investigation of the orbit structure of this action.